OpenStudy (anonymous):

A data set consists of the following data points: (2,12), (3,7), (5,4) The slope of the best fit line is -2.5. Find the y-intercept of this line.

OpenStudy (anonymous):

@amistre64 @skinny23

OpenStudy (anonymous):

@beutygirl658

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

y - y₁ = m(x-x₁) y - 12 = -2.5(x - 2) y - 12 = -2.5x + 5 y = -2.5x + 17 The y-intercept is 17.

OpenStudy (anonymous):

It said the answer was 16. Oh well, thank you for your help!!! :)